# INTRABENCH: Interactive Radiological Benchmark

### v1

@@ Abstract
Current interactive segmentation approaches, inspired by the success of META’s Segment Anything model, have achieved notable advancements, however they come with substantial limitations that hinder their practical application in real clinical scenarios. These include unrealistic human interaction requirements, such as slice-by-slice operations for 2D models on 3D data, a lack of iterative refinement, and insufficient evaluation experiments. These shortcomings prevent accurate assessment of model performance and lead to inconsistent outcomes across studies.
IntRaBench  overcomes these challenges by offering a comprehensive and reproducible framework for evaluating interactive segmentation methods in realistic, clinically relevant scenarios. It includes diverse datasets, target structures, and segmentation models, and provides a flexible codebase that allows seamless integration of new models and prompting strategies. Additionally, we introduce advanced techniques to minimize clinician interaction, ensuring fair comparisons between 2D and 3D models. By open-sourcing IntRaBench, we invite the research community to integrate their models and prompting techniques, ensuring continuous and transparent evaluation of interactive segmentation models in 3D medical imaging.
@@ 1 Introduction
Accurate segmentation of anatomical structures or pathological areas is crucial in fields like radiology, oncology, and surgery to isolate affected regions, monitor disease progression, treatment planning and guide therapeutic procedures. Traditional supervised medical segmentation models have demonstrated strong performance across a range of anatomies and pathologies (Isensee et al., 2020; 2023; Huang et al., 2023; Ulrich et al., 2023). However, their effectiveness remains heavily constrained by the amount and diversity of available training data, with the quality of human label annotations serving as a critical limiting factor. Consequently, fully autonomous AI solutions have not yet reached performance needed for widespread autonomous clinical applications.
@@
On the other hand, numerous semi-automatic segmentation techniques, not reliant on AI, are already in clinical practice to expedite manual annotation processes Hemalatha et al. (2018). These current ad hoc methods do not tap into the potential of AI-based automation to drastically reduce annotation time. A method that allows clinicians to segment any target with just a single click within the image could greatly enhance the efficiency of clinical workflows.
@@
The release of META’s Segment Anything (SAM) model represents a big leap towards making this potential a reality (Kirillov et al., 2023). ”SAM” is designed to segment any target through different user interaction methods, including point-based and bounding box prompts. This allows users to easily specify the area of interest by clicking on it or drawing a bounding box around it, making the segmentation process both flexible and intuitive. A particularly powerful feature is the ability for users to iteratively refine initial predictions by adding more positive or negative prompts.
@@
This advanced functionality, in contrast to traditional supervised segmentation methods, has attracted a lot of attention in the medical domain, and led to many studies evaluating and adapting SAM for 3D medical image segmentation (Roy et al., 2023; Deng et al., 2023; Hu et al., 2023; Zhou et al., 2023; Mohapatra et al., 2023; Cheng et al., 2023; Ma et al., 2024; Gong et al., 2023). Moreover, several researchers have been inspired by SAM’s capabilities to develop their own methods, often specifically designed for the 3D nature of radiological data (Du et al., 2024; He et al., 2024; Li et al., 2024; Wang et al., 2024).
@@
Although these domain-specific adaptations on medical data have shown promising progress, many published methods are plagued by pitfalls which obfuscate the efficacy of the models and prevent clinicians and researchers from determining the best methods for their use-cases:
@@
Figure 1: a) Current approaches require clinicians to interact with radiological images slice by slice, leading to increased workload. b) Some models operate natively in 3D and enable full 3D interaction. Only models that accept mask prompts allow iterative refinement of initial predictions with human guidance.
@@
Applying interactive 2D models to 3D data on a slice-by-slice basis (P1): Assuming clinicians will interact with each slice individually is unrealistic and undermines the efficiency improvements these methods aim for. Moreover, a slice-by-slice approach introduces an unfair bias when comparing 2D and 3D models, as 3D models typically require only a few interaction per image, leading to significantly fewer interactions and less supervision Cheng et al. (2023); Ma et al. (2024); Zhang & Liu (2023); Wu et al. (2024); Wong et al. (2024).
@@
Neglecting refinement (P2): Many studies assess interactive segmentation methods based on a single interaction step, overlooking the inherent ambiguities in radiological images (Ma et al., 2024; Du et al., 2024; Gong et al., 2023; Bui et al., 2024). Often, a second interaction may be necessary to specify which specific substructure the clinician wants to segment. This could be, e.g. a vessel within the liver, or the necrosis within a tumor, as exemplified in the well-known BraTs segmentation challenge (de Verdier et al., 2024). Furthermore, clinicians often want to adapt the segmentations to their clinic’s local protocol or refine them, particularly for targets with high inter-rater variability, like pathological structures (Fu et al., 2014; Benchoufi et al., 2020; Hesamian et al., 2019). Overall, there is a notable lack of research exploring realistic, iterative interaction methods for 2D models applied to 3D volumes.
@@
Obfuscated and insufficient evaluation (P3): With promptable models only recently garnering great attention, there is a lack of a standardized approach to evaluation, which has led to disparate and incomparable methods, which are at times even obfuscated or insufficient. We observed the following shortcomings: (i) Not specifying whether predictions were interactively refined or based on a single prompt with multiple points (Cheng et al., 2023; Wang et al., 2024). (ii) Being intransparent on the number of initial prompts given (Du et al., 2024). (iii) Using the best mask rather than the final mask after interactive refinement (Wang et al., 2024). (iv) Evaluating predictions slice-by-slice or on sub-patches of a 3D volume instead of evaluating on the full image (Roy et al., 2023; Ma et al., 2024; Cheng et al., 2023; He et al., 2024; Li et al., 2024). (v) Excluding targets considered ’too small’, hence neglecting valid targets such as small lesions that are neither tested nor trained on Ma et al. (2024); Cheng et al. (2023); Wang et al. (2024). (vi) Comparing against non-promptable models and SAM, rather than any other promptable model trained on medical data (Cheng et al., 2023; Ma et al., 2024; Gong et al., 2023; He et al., 2024). (vii) Lastly, overemphasizing segmenting healthy structures, such as organs, where existing supervised public models already perform well (Wasserthal et al., 2023; Ulrich et al., 2023), instead of focusing on pathologies, where interactive refinement could provide the greatest benefits (Wang et al., 2024; Zhang & Liu, 2023).
To address these pitfalls, a benchmark is needed, aligning with the recent review paper from Marinov et al. (2024). To this end, we introduce IntRaBench, a reproducible and extendable Interactive Radiological Benchmark. Through it, we highlight the most performant 2D and 3D interactive segmentation and the best prompting methods in the radiological domain. In this paper, we present experiments carefully designed to replicate a clinical workflow as closely as possible, with the following key contributions:
@@
    IntRaBench, for the first time, enables a fair comparison of the most influential 2D and 3D interactive segmentation methods. By measuring the number of simulated interactions, a proxy for the ’Human Effort’, we test different prompting strategies that do not require a slice-wise interaction (P1).
@@
    The extendable IntRaBench framework allows developers to a) easily evaluate a new method in a fair manner against established methods and b) easily develop and investigate new prompting strategies.
Through open-sourcing IntRaBench, we invite researchers to integrate their methods into our framework, promoting continuous and equitable assessment that allows us to track the overall progress in the field of interactive 3D medical image segmentation reproducibly and transparently.
Figure 2: IntRaBench overview. Although our evaluation is performed on entire 3D volumes, the benchmark accommodates both 3D and 2D interactive segmentation methods. While 3D model prompting is relatively straightforward, we introduce prompting and refinement strategies for 2D models that minimize the effort required from human interaction. The benchmark is designed to be extensible, and researchers are encouraged to propose and integrate additional methods seamlessly using our codebase particularly for areas marked by three dots.
2 IntRaBench
The Interactive Radiology Benchmark is designed to easily enable a fair and reproducible evaluation of 2D and 3D interactive segmentation methods for 3D radiological image segmentation for the very first time. While prompting 3D models is generally straightforward, we introduce specific prompting and refinement strategies for 2D models to streamline human interaction and reduce the simulated ’Human Effort’. The proposed benchmark includes seven established models and ten datasets covering different target structures and image modalities. All datasets are publicly available and we support an automatic download and preprocessing for improved usability and reproducibility.
Moreover, the benchmark is built with flexibility in mind, enabling seamless integration of additional methods, as visualized in Fig. 2. Researchers are invited to contribute new approaches, particularly new models, new prompting schemes, and new interesting datasets to the collection. Overall, the design of our benchmark allows for easy testing and validation of novel segmentation methods, making the benchmark a catalyst for advancing methodology for interactive 3D medical image segmentation. In the following, we present the different components of IntRaBench.
@@ 2.1 Initial Prompting
Prompts are a key component of any interactive segmentation method and can highly influence the overall performance of the underlying method. IntRaBenchdistinguishes between two visual prompting types. Point prompts correspond to a click of a user in the image, and box prompts refer to a box around the target structure. While there is no difference in providing a point prompt for 2D and 3D methods, a 3D box requires an additional dimension compared to a 2D box. Notably, some methods also enable a distinction between foreground and background point prompts. While 3D models allow segmenting a 3D volume natively, 2D-based models require an interaction for each slice, resulting in excessive effort, which is prohibitive for clinicians as it would take too much time in daily clinical practice. Hence, any meaningful performance comparison must account for this difference in prompting effort.
To increase the feasibility of 2D models for 3D applications, it is essential to reduce this effort. We propose two straightforward methods, for both point and box prompts, to explore their performance and provide a proxy for measuring the effort of human interaction.
Point interpolation: Let $I\subset N$ be a set of axial indices of all foreground slices. We simulate a user by selecting $n$ foreground points, specifically the center of the largest connected component of slice $i_{1},...i_{n}\in I$ where the $i_{j}$ are equally spaced within $I$ and $i_{1}=\min(I)$ and $i_{n}=\max(I)$. Then, we interpolate linearly between each point and the next one and use the intersections of the resulting lines with the axial slices as positive point prompts, as visualized in Fig. 3 c).
Point propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D point prompt within the median slice corresponding to the median axial index $i_{m}$. Given this point, the model generates a segmentation $S_{m}$ for the median slice. We then calculate a ’central point,’ specifically the center of mass of the largest connected component of $S_{m}$, to use as a point prompt for the slice indexed by $i_{m-1}$. Again, we generate a segmentation $S_{m-1}$ of this slice, and create a new central point until we segment the slice with the axial index $\min(I)$. The propagation is then repeated upwards, starting from $i_{m+1}$ and continuing until we segment the slice with the axial index $\max(I)$. This process is visualized in Fig. 3 e).
@@
Box propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D box prompt within the slice corresponding to the axial index $i_{m}$, $m$ as in point propagation. The model then generates a segmentation $S_{m}$ for the median slice. A new bounding box is created based on $S_{m}$ and used as a prompt for the slice indexed by $i_{m}-1$. The propagation is continued down to $\min(I)$ and then repeated upwards until $\max(I)$ as in point propagation, but using box prompts instead of point prompts. See Fig. 3 f) for a visualization.
While one cares about realistic prompting behavior, IntRaBench also supports the previously mentioned slice-by-slice prompting styles for completeness.
@@ 2.2 Refinement Prompting
Refinement of previous segmentations is an important aspect of interactive segmentation models, as it allows iteratively improving the segmentation until the desired structure is segmented to a user’s demands. Some interactive segmentation models allow for the refinement of initial segmentations by providing the model with the previous prediction along with a new prompt to correct errors, either through foreground clicks on false negative pixels or background clicks on false positive pixels. While this process is straightforward for 3D models, 2D models naively only allow for refinement in a slice-by-slice fashion, which again places an unrealistic burden on clinicians. Therefore, we present refinement strategies that require a reasonable level of ”Human Effort”.
@@
If positive prompts are selected, we perform a connected component analysis on the false negative voxels. Given $L$, the largest connected component, we generate a scribble from the bottom to the top of $L$ by taking the centroid of $L$ in each slice to simulate drawing a vertical scribble through the ’middle’ of $L$. This simulates a clinician annotating regions that were erroneously not segmented.
@@
If negative prompts are selected, we identify a non-axial slice $S_{fp}$ of $S$ that contains the most false positives. Then we generate a contour curve around the ground truth target object at a distance of 2 pixels. We then select a subpart $C$ with a length of 60% of the full curve and sample all pixels $c\in C$ that are false positives to obtain a set of points $D$, simulating a user drawing a few scribbles in areas where the model over-segmented the target. For 2D models, we then generate new slice predictions for each slice containing a point in $D$ by providing the model with the previous prediction as well as new negative prompts: all $d\in D$ which belong to that slice. For 3D models, we again feed the whole 3D patch, together with the previous prediction $S$ and a negative prompt sampled from $D$.

### v2

# RadioActive: 3D Radiological Interactive Segmentation Benchmark

@@ Abstract
Current interactive segmentation approaches, inspired by the success of META’s Segment Anything model, have achieved notable advancements, however, they come with substantial limitations that hinder their practical application in 3D radiological scenarios. These include unrealistic human interaction requirements, such as slice-by-slice operations for 2D models on 3D data, a lack of iterative interactive refinement, and insufficient evaluation experiments. These shortcomings prevent accurate assessment of model performance and lead to inconsistent outcomes across studies. The RadioActive benchmark overcomes these challenges by offering a comprehensive and reproducible evaluation of interactive segmentation methods in realistic, clinically relevant scenarios. It includes diverse datasets, target structures, and interactive segmentation methods, and provides a flexible, extendable codebase that allows seamless integration of new models and prompting strategies. We also introduce advanced prompting techniques to enable 2D models on 3D data by reducing the needed number of interaction steps, enabling a fair comparison. We show that surprisingly the performance of slice-wise prompted approaches can match native 3D methods, despite the domain gap. Our findings challenge the current literature and highlight that models not specifically trained on medical data can outperform the current specialized medical methods. By open-sourcing RadioActive, we invite the research community to integrate their models and prompting techniques, ensuring continuous and transparent evaluation of interactive segmentation models in 3D medical imaging.
@@ 1 Introduction
Figure 1: Current interactive segmentation methods require clinicians to interact with radiological images slice-by-slice, leading to increased workload.
@@
Accurate segmentation of anatomical structures or pathological areas is crucial in fields like radiology, oncology, and surgery to isolate affected regions, monitor disease progression, and treatment planning, and guide therapeutic procedures. On the one hand, traditional supervised medical segmentation models have demonstrated strong performance across a range of anatomies and pathologies [18, 19, 17, 37]. However, their applicability remains heavily constrained by the limited task-specific training data available, as well as their deterministic predictive behavior that cannot be influenced, possibly requiring labor-intensive manual post-processing of predictions. Consequently, fully autonomous AI solutions have not yet found widespread autonomous clinical applications.
@@
On the other hand, numerous semi-automatic segmentation techniques, not reliant on AI, are already in clinical practice to expedite manual annotation processes [13]. These current ad hoc methods do not yet tap into the potential of AI-based automation to drastically reduce annotation time. A method that bridges this gap and allows clinicians to segment any target with just a single click within the image could greatly enhance the efficiency of clinical workflows.
@@
Interactive segmentation methods, like the Segment Anything Model (SAM) [20] represent exactly this. SAM is designed to segment any target through various user interaction methods, including point-based and bounding box prompts. This allows users to easily specify the area of interest by clicking on it or drawing a bounding box around it, making the segmentation process both flexible and intuitive. A particularly powerful feature is the ability for users to iteratively refine initial predictions by adding more positive or negative prompts.
@@
Due to their great potential, interactive segmentation methods have attracted a lot of attention in the medical domain and led to many studies evaluating and adapting SAM for 3D medical image segmentation [34, 5, 15, 45, 30, 3, 24, 11]. Moreover, several researchers have been inspired by SAM’s capabilities to develop their own methods, often specifically designed for the 3D nature of radiological data [9, 12, 21, 39].
@@
Although these domain-specific adaptations to medical data have shown promising progress, many published methods are plagued by pitfalls that obfuscate the efficacy of the models and prevent clinicians and researchers from determining the best methods for their use cases:
@@
Figure 2: Some models operate natively in 3D and enable full 3D interaction. Only models that accept mask prompts allow iterative refinement of initial predictions with human guidance.
@@
Applying interactive 2D models to 3D data on a slice-by-slice basis (P1): Assuming clinicians will interact with each slice individually is unrealistic and undermines the efficiency improvements these methods aim for. Moreover, a slice-by-slice approach introduces an unfair bias when comparing 2D and 3D models, as 3D models typically require only a few interactions per image, leading to significantly fewer interactions and less supervision [3, 24, 43, 42, 41, 27].
@@
Neglecting refinement (P2): Many studies assess interactive segmentation methods based on a single interaction step, overlooking the inherent ambiguities in radiological images [24, 9, 11, 2]. Often, a second interaction may be necessary to specify which specific substructure the clinician wants to segment. This could be, e.g. a vessel within the liver, or the necrosis within a tumor, as exemplified in the well-known BraTs segmentation challenge [4]. Furthermore, clinicians often want to adapt the segmentations to their clinic’s local protocol or refine them, particularly for targets with high inter-rater variability, like pathological structures [10, 1, 14]. Overall, there is a notable lack of research exploring realistic, iterative interaction methods for 2D models applied to 3D volumes.
@@
Obfuscated and insufficient evaluation (P3): Currently, there is a lack of a standardized approach to evaluation. Often results are not incomparable between papers, with evaluation often obfuscated or insufficient. We observed the following shortcomings: (i) Not specifying whether predictions were interactively refined or based on a single prompt with multiple points [3, 39]. (ii) Being intransparent on the number of initial prompts given [9]. (iii) Using the best mask rather than the final mask after interactive refinement [39]. (iv) Evaluating predictions slice-by-slice or on sub-patches of a 3D volume instead of evaluating on the full image [34, 24, 3, 12, 21]. (v) Excluding targets considered too small, hence neglecting valid targets such as small lesions that are neither tested nor trained on [24, 3, 39]. (vi) Comparing against non-promptable models and SAM, rather than any other promptable model trained on medical data [3, 24, 11, 12]. (vii) Lastly, overemphasizing segmenting healthy structures, such as organs, where existing supervised public models already perform well [40, 37], instead of focusing on pathologies, where interactive refinement could provide the greatest benefits [39, 43].
To overcome these challenges, establishing a benchmark is essential, aligning with the recent recommendations by Marinov et al. [26]. To this end, we present RadioActive, a robust, reproducible, and extendable Radiological InterActive Benchmark for 3D medical image segmentation. This benchmark highlights the most effective 2D and 3D interactive segmentation approaches and tests realistic prompting methods in the 3D domain. We conduct evaluations directly in the 3D domain and across multiple models, providing more generalizing insights, as opposed to prior work [16, 26, 34]. Our key contributions include:
@@
    RadioActive, for the first time, enables a fair comparison of the most influential 2D and 3D interactive segmentation methods for 3D medical data. By measuring the number of simulated interactions, a proxy for the ’Human Effort’, we test different prompting strategies that do not require slice-by-slice interactions (P1).
@@
    The extendable RadioActive framework allows developers to a) easily evaluate a new method in a fair manner against established methods and b) easily develop and investigate new prompting strategies.
Through open-sourcing RadioActive, we invite researchers to integrate their methods into our framework, promoting continuous and equitable assessment that allows us to track the overall progress in the field of interactive 3D medical image segmentation reproducibly and transparently.
Figure 3: RadioActive overview. Although our evaluation is performed on entire 3D volumes, the benchmark accommodates both 3D and 2D interactive segmentation methods. While 3D model prompting is relatively straightforward, we introduce prompting and refinement strategies for 2D models that minimize the effort required from human interaction. The benchmark is designed to be extensible, and researchers are encouraged to propose and integrate additional methods seamlessly using our codebase particularly for areas marked by three dots.
2 RadioActive
The RadioActive Benchmark is designed to easily enable a fair and reproducible evaluation of 2D and 3D interactive segmentation methods for 3D radiological image segmentation for the very first time Fig. 3. While prompting 3D models is generally straightforward, we introduce specific prompting and refinement strategies for 2D models to streamline human interaction and reduce the simulated Human Effort. The proposed benchmark includes seven established models and ten datasets covering different target structures and image modalities. All datasets are publicly available and we provide an automatic download and pre-processing for improved usability and reproducibility.
Moreover, the benchmark is built with flexibility in mind, enabling seamless integration of additional methods. Researchers are invited to contribute new approaches, particularly new models, new prompting schemes, and new interesting datasets to the collection. Overall, the design of our benchmark allows for easy testing and validation of novel segmentation methods, making the benchmark a catalyst for advancing methodology for interactive 3D medical image segmentation. In the following, we present the different components of RadioActive.
@@ 2.1 Initial Prompting
Prompts are a key component of any interactive segmentation method and can highly influence the achieved performance of the underlying method. RadioActive distinguishes between two visual prompting types. Point prompts correspond to a click of a user in the image, and box prompts refer to a box around the target structure. Notably, some methods also enable a distinction between foreground and background point prompts. While 3D models allow segmenting a 3D volume natively, 2D-based models require an interaction for each slice, resulting in excessive effort, which is prohibitive for clinicians as it would take too much time in daily clinical practice. Hence, any meaningful benchmark must account for this difference in prompting effort. To increase the feasibility of 2D models for 3D applications, it is essential to reduce this effort. We propose two straightforward methods, for both point and box prompts, to explore their performance and provide a proxy for measuring the effort of human interaction.
Point interpolation: Let $I\subseteq N$ be the set of axial indices of all foreground slices. We simulate a user by selecting $n$ foreground points, specifically the center of the largest connected component of slice $i_{1},...,i_{n}\in I$ where the $i_{j}$ are equally spaced within $I$ and $i_{1}=\min(I)$ and $i_{n}=\max(I)$. Then, we interpolate linearly between each point and the next one and use the intersections of the resulting lines with the axial slices as positive point prompts, as visualized in Fig. 4 c).
Point propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D point prompt within the slice corresponding to the median axial index $i_{m}$. Given this point, the model generates a segmentation $S_{m}$ for the median slice. We then calculate a central point, specifically the center of mass of the largest connected component of $S_{m}$, to use as a point prompt for the slice indexed by $i_{m-1}$. Again, we generate a segmentation $S_{m-1}$ of this slice, and create a new central point until we segment the slice with the axial index $\min(I)$. The propagation is then repeated upwards, starting from $i_{m+1}$ and continuing until we segment the slice with the axial index $\max(I)$. This process is visualized in Fig. 4 e).
@@
Box propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D box prompt within the slice corresponding to the axial index $i_{m}$, with $m$ defined as in point propagation. The model then generates a segmentation $S_{m}$ for the median slice. A new bounding box is created based on $S_{m}$ and used as a prompt for the slice indexed by $i_{m}-1$. The propagation is continued down to $\min(I)$ and then repeated upwards until $\max(I)$ as in point propagation, but using box prompts instead of point prompts. See Fig. 4 f) for a visualization.
While realistic prompting behavior is the priority RadioActive also supports the previously used unrealistic slice-by-slice prompting styles for completeness. Additional details on prompt schemes are provided in the appendix in Sec. 6.
@@ 2.2 Refinement Prompting
Refinement of previous segmentations is an important aspect of interactive segmentation models, as it allows iteratively improving the segmentation until the desired structure is segmented to a user’s demands. Some interactive segmentation models allow for the refinement of initial segmentations by providing the model with the previous prediction along with a new prompt to correct errors, either through foreground clicks on false negative pixels or background clicks on false positive pixels. While this process is straightforward for 3D models, 2D models naively only allow for refinement in a slice-by-slice fashion, which again places an unrealistic burden on clinicians. Therefore, we present refinement strategies that require a reasonable level of Human Effort.
@@
If positive prompts are selected, we perform a connected component analysis on the false negative voxels. Given $L$, the largest connected component, we generate a scribble from the bottom to the top of $L$ by taking the centroid of $L$ in each slice to simulate drawing a vertical scribble through the middle of $L$. This simulates a clinician annotating regions that were erroneously not segmented.
@@
If negative prompts are selected, we identify a non-axial slice $S_{fp}$ of $S$ that contains the most false positives. Then we generate a contour curve around the ground truth target object at a distance of 2 pixels. We then select a subpart $C$ with a length of 60% of the full curve and sample all pixels $c\in C$ that are false positives to obtain a set of points $D$, simulating a user drawing a few scribbles in areas where the model over-segmented the target. For 2D models, we then generate new slice predictions for each slice containing a point in $D$ by providing the model with the previous prediction as well as new negative prompts: all $d\in D$ which belong to that slice. For 3D models, we again feed the whole 3D patch, together with the previous prediction $S$ and a negative prompt sampled from $D$. The motivation for the different approaches for FPs is, that correcting the model near the target boundary is likely to provide more informative feedback. As none of the evaluated methods features a dedicated scribble prompt encoding, we instead generate a set of point prompts from our scribble refinement trajectory, which is similar to the recently proposed work of Wong et al. [41] for scribble prompting of SAM. Aside from this refinement strategy, we also support unrealistic refinement prompts for which details are provided in the appendix in Sec. 6.